### v1

@@ 3.2 Contrast
is the resist thickness, denotes the dose, is the layer thickness for full exposure, and is a parameter that accounts for the exponential’s slope and consequently the contrast . To find , we calculate where the horizontal line intersects the tangent to the fit function at . Fig. 2 (a) shows the measured contrast curves. For better comparison, we normalize both contrast curves to their respective maxima. On average, using the spin coating recipe described above, we achieve with Medusa 84 SiH a layer thickness of = 150.7 nm 0.3 nm on silicon at a maximum rotational speed of 4000 rpm. The initial 60 m squares emerge above = 55.0 C/cm2 0.5 C/cm2, and we identified a base dose of = 104.1 C/cm2 0.9 C/cm2. We find = 3.6 0.08, similar to FOx 16 = 4.5 0.2 under similar processing conditions in our previous work [20, 21, 6, 7]. We note that the FOx 16 layer was almost 700 nm thicker and we used a different developer so the results are only partly comparable.
@@ 3.3 Selectivity
Furthermore, we examine if the etch resistance of Medusa 84 SiH depends on the dose. We expose our above discussed contrast curve pattern for 1 minute to the ”Pillar Etch” plasma and a contrast curves pattern on a second silicon substrate for 45 s to the biased SF6 plasma used for silicon adhesion layer removal. To determine the selectivity, we use the following approach. We first measure the surface profile of the sample after ICP-RIE via profilometry. We then remove the Medusa 84 SiH mask using BOE treatment which induces negligible etching of the silicon substrate. We again measure the surface profile of the silicon substrate. We now extract the resist thickness via comparing the surface profiles to account for etching of the silicon substrate in the plasma. Utilizing the resist’s thicknesses before and after ICP-RIE, we compute the etch rates of the plasmas for the materials involved and determine selectivity ratios. Fig. 2 (b),(c) summarizes the resulting, dose dependent selectivities (details see caption). For comparison, we repeat this process on each ICP-RIE etched diamond using three randomly selected 25 m square structures. On diamond, we observe a decreased etchrate of the ”Pillar Etch” plasma on Medusa 84 SiH compared to silicon leading to an increased average selectivity of 11 6 for the chosen dose of 350 C/cm2. This indicates that the insulating diamond substrate influences plasma dynamics and reduces the etch rate of the ”Pillar Etch” on Medusa 84 SiH by approximately 30%, favoring its application in diamond nanofabrication. The selectivity values presented in this study are comparable to those previously reported for similar ICP-RIE recipes using HSQ or SiO2-like masks [25, 4].
@@ 3.4 Fabrication Yield
We analyze optical microscopy images of all diamond nanopillars fields to quantify the fabrication yield of nanopillars. We note that prior to this evaluation, we used SEM imaging to make sure the structures we observe in optical microscopy are intact nanopillars and not structures created due to collapsed pillar masks that adhered to their original positions. A total of 79.200 nanopillars have been evaluated on a single electronic grade diamond. We report a total fabrication yield of 89.6%. Among the 10.4% fabrication failures, 59.4% result from a non-optimal synthesis and oligomeric silsesquioxanes, 20.7% are caused by edge beads that affect lithography and 17.7% are the result of adhesion failure. The remaining 2.2% is due to contaminants, as well as overexposure in lithography. The work presented here is based on an early prototype of Medusa 84 SiH by Allresist. The now commercially available version has been significantly optimized regarding manufacturing to remove the effect originating from non-optimal synthesis. Therefore fabrication yields could potentially reach around 98%, primarily constrained by edge beads on small diamond substrates and sporadic adhesion failures.
@@ 4 Experimental Setup
We used a custom-built confocal microscope (modified version of [26]) to study the impact of diamond nanostructuring with Medusa 84 SiH on NV center spin properties. We utilize a diode laser at 520 nm, featuring sub-nanosecond digital modulation (DL nSec, PE 520, Swabian Instruments) and a maximum output power of 40 mW to excite and initialize the NV centers. A 600 nm shortpass filter placed behind the single-mode fiber guiding the excitation light eliminates background light from the fiber. The microscope objective (LMPLFLN100X, Olympus) has a numerical aperture of 0.8, along with a working distance of 3.4 mm. We utilize a piezo-electric scanner (ANSxyz100, standard range, Attocube Systems) to facilitate sample scanning. We use two dichroic mirrors and a 600 nm longpass filter to separate excitation and fluorescence light. We sent the fluorescence via a single mode fiber to a fiber coupled avalanche photo diode (APD, SPCM-AQRH-14-FC, Excelitas Technologies) with a quantum efficiency of approximately 69 % at 700 nm. A data acquisition card (PCIe-6323, National Instruments) acquires and logs the signal. To analyze fluorescence spectra, we direct the light to a diffraction grating spectrometer (Acton SpectraPro SP-2500, Princeton Instruments) that records the spectra using a CCD camera (Pixis 256E, Princeton Instruments) using a multi-mode fiber replacing the single mode detection fiber. To investigate excited state lifetimes via time correlated single photon counting, we use a supercontinuum laser (SuperK Extreme EXW-12, NKT-Photonics) with a configurable single-line filter (SuperK VARIA, NKT-Photonics) and a time-tagging module (PicoHarp 300, PicoQuant). For coherent spin manipulation, a microwave source (TSG4100A, Tektronix) is connected to a +45 dB amplifier (ZHL-16W-43+, Mini-Circuits), feeding a -shaped stripline microwave antenna [27], on which the diamond is fixed using Crystal Bond 509 (Structure Probe Inc). We apply adjustable external magnetic fields using neodymium magnets on a rotation mount and a three-axis linear translation stage.
@@ 5 Shallow NV Centers in Nanopillars manufactured using Medusa 84 SiH Masks
Fig. 3 (a) illustrates a confocal xy-scan of a nanopillar array (nominal diameters 180 nm). The fluorescence intensity emitted from the nanopillars is typical for individual NV centers in such structures [30, 6, 7, 5, 4]. Fluorescence spectra of pillars reveal the spectrum of NV centers, characterized by their broad phononic sideband and a zero phonon line at 637 nm, as shown in the inset of Fig. 3 (b). We perform lifetime measurements and fit them using a bi-exponential decay convoluted with a Gaussian to account for the instrument response function of our setup (244 ps width) as shown in Fig. 3 (b). Using the bi-exponential decay is necessary as we observe a fast decaying fluorescence component that we attribute to (broadband) surface fluorescence ( 1.9 ns 0.1 ns) [28, 29]. We determine the NV excited state lifetime to be 15 ns 2 ns, which is slightly increased in comparison to NV centers in bulk diamond due to their proximity to the surface. To estimate the number of NV centers per nanopillar, we evaluated xy-scans for nanopillar fields. By identifying the local maxima of the histogram in Fig. 3 (c) as occurring due to the presence of 0, 1, 2, or 3 NV centers and assuming Gaussian distributions, we approximated the occurrence of a single NV center per nanopillar as 18 % 7 % (details see caption). Using this distribution, we estimate the creation yield of shallow NV centers in our sample. We determined an NV center density of , indicating a creation yield of which is comparable to other works [6]. We note that this approach only roughly estimates the NV density as the expected photonic enhancement and thus the apparent brightness of NV centers in nanopillars has been found to strongly depend on the spatial position within the nanopillar (see, e.g. Ref. [7]). We note that the observed NV centers are stable and do not show blinking or bleaching. This observation together with the NV creation yield matching previous work indicate that Medusa 84 SiH efficiently protects shallow NV centers during ICP-RIE.
@@
We examine the spin characteristics of individual NV centers in the fabricated nanopillars to check if these shallow NV centers would be suitable as quantum sensors. To this end, we perform optically detected magnetic resonance (ODMR) measurements in presence of an external magnetic field and select nanopillars for further investigation which show only a single resonance pair and therefore with high probability contain only a single NV center. We select a single transition for resonant driving of Rabi oscillations, to determine the -pulse duration as shown in the inset of Fig. 3 (d). Subsequently, we perform spin echo experiments to determine the NV center’s coherence time , as illustrated in Fig. 4 (a). This procedure was repeated for 17 nanopillars, each containing a single NV center, to compare with reference values determined on the same diamond after NV creation followed by Tri-Acid cleaning and before proceeding with processing using Medusa 84 SiH. We employ a fitting function from [26] to extract :
@@
For the NV centers investigated in this study, we found an average coherence time of = 14 s 9 s before nanopillar fabrication using Medusa 84 SiH and = 9 s 6 s after processing. As shown in Fig. 4 (b), the coherence times are in the same window as observed before, which indicates that processing diamond using Medusa 84 SiH does not have a considerable effect on shallow NV center’s coherence times T2. Determined coherence times fall into the same order of magnitude as reported for single NV centers within nanopillars in the literature [6]. Additionally, Ramsey measurements, as displayed in Fig. 4 (c), have been conducted. In average, we found dephasing times of = 1.2 s 0.4 s which lies in the same range as for commercially available single NV scanning probes [8]. Moreover, T1 spin lifetime measurements were performed before and after processing. As shown in Fig. 4 (d), the spin lifetime also shows only a slight decrease after processing (average T1 = 0.3 ms 0.1 ms ). While T1 before and after processing is consistent for our diamond sample, we note that we typically find T1 times of several milliseconds for commercially available NV scanning probes [8]. These results conclusively demonstrate that the use of Medusa 84 SiH for diamond nanofabrication, compared to the cutting-edge HSQ resist FOx by DuPont, does not adversely affect the spin properties of color centers, thus rendering Medusa 84 SiH appropriate for diamond-based quantum technologies.
@@ 6 Conclusion
This study examined Allresist’s novel electron beam resist Medusa 84 SiH as a substitute for DuPont’s Flowable Oxide for diamond nanofabrication, focusing on its use in diamond-based quantum technologies. We transferred previously established processes for FOx to Medusa 84 SiH [6, 7, 21] and proved their feasibility by realizing prototypical test structures at the m-scale down to nanopillars with diameters around 200 nm. We investigated the resist’s basic properties by evaluating fabrication yield, recording contrast curves and determining its dose-dependent etch resistance against common plasmas necessary to nanostructure diamond. We confirm that Medusa 84 SiH is very similar to FOx with comparable contrast, achieved fabrication yields of up to 98% for electron beam evaporated silicon as adhesion promoting interlayer and shows a minimum selectivity towards high biased O2 + Ar ICP-RIE plasma of 6 1 for m-scale structures. We further investigated the impact of processing diamond with Medusa 84 SiH on shallow NV centers within nanopillars as applied in the field of quantum sensing [6, 7, 8] by comparing T2, T and T1 times before and after fabricating nanopillars. We report negligible decreases in spin lifetime, dephasing and coherence times, therefore proving the suitability of these nanopillars for quantum measurement schemes. Further, we investigated their fluorescence emission and show that, with the proposed nanofabrication process, an enhancement in photon collection efficiency of 7.0 0.9 was achieved. Overall, Medusa 84 SiH is a promising alternative to FOx as an electron beam resist for diamond nanofabrication compatible with quantum technologies like NV sensing. Furthermore, this work leaves room for further optimization, as the resist’s contrast is expected to achieve even higher values by lowering the exposure to thermal energy during soft baking and adjusting the development parameters. In addition to diamond, this work hints Medusa 84 SiH might be also an appropriate alternative for FOx in other material systems like silicon, SOI, GaAs or other materials under research requiring nanoscale resolution and high resistance towards etch processes.
@@ 7.1 Nanofabrication
| Spin Coating of Resist Layer | Step #1: 1500 rpm for 6 s, Acceleration: 500 rpm/s |
| Step #2: 4000 rpm for 30 s, Acceleration: 1000 rpm/s | |
@@
| Application of conductive Layer | 4000 rpm for 30 s, Acceleration: 1000 rpm/s |
| Resist: ESpacer 300Z | |
@@
| Rinse: Ultrapure water for 30 s | |
| Cleaning: Isopropanol for 30 s | |
@@
Table 1 provides a comprehensive summary of all process steps in sequential order, detailing the majority of the process parameters specifically related to the electronic grade diamond. For diamonds of lower purity utilized in process development, the stress relief etch, ion implantation, and ultra high vacuum annealing steps were omitted. Moreover, the silicon samples used for contrast curve analysis did not undergo tri-acid or KOH treatments, and procedures such as electron beam evaporation, deposition of a conductive layer, and proximity correction were also excluded.
@@
Table 2 summarizes the parameters used to establish the ICP-RIE plasmas used to conduct stress relief etches on an Oxford Instruments Plasmalab 100 or to structure a diamond with the proposed fabrication process using a Sentech PTSA-ICP Plasma Etcher SI 500.

### v2

@@ 3.2 Contrast
is the resist thickness, denotes the dose, is the layer thickness for full exposure, and is a parameter that accounts for the exponential’s slope and consequently the contrast . To find , we calculate where the horizontal line intersects the tangent to the fit function at . Fig. 2 (a) shows the measured contrast curves. For better comparison, we normalize both contrast curves to their respective maxima. On average, using the spin coating recipe described above, we achieve with Medusa 84 SiH a layer thickness of = 150.7 nm 0.3 nm on silicon at a maximum rotational speed of 4000 rpm. The initial 60 m squares emerge above = 55.0 C/cm2 0.5 C/cm2, and we identified a base dose of = 104.1 C/cm2 0.9 C/cm2. We find = 3.6 0.08, similar to FOx 16 = 4.5 0.2 under similar processing conditions in our previous work [20, 21, 6, 7]. We note that the FOx 16 layer was almost 700 nm thicker and we used a different developer so the results are only partially comparable.
@@ 3.3 Selectivity
Furthermore, we examine if the etch resistance of Medusa 84 SiH depends on the dose. We note that for thick resist layers, the selectivity might vary with thickness because of depth-dependent effective dose distribution within the layer. Consequently, thickness variations might influence the evaluation of selectivity especially for diamond where thickness varies stronger. We expose our above discussed contrast curve pattern for 1 minute to the ”Pillar Etch” plasma and another pattern on a second silicon substrate for 45 s to the biased SF6 plasma used for silicon adhesion layer removal. To determine the selectivity, we use the following approach. We first measure the surface profile of the sample after ICP-RIE via profilometry. We then remove the Medusa 84 SiH mask using BOE treatment which induces negligible etching of the silicon substrate. We again measure the surface profile of the silicon substrate. We now extract the resist thickness via comparing the surface profiles to account for etching of the silicon substrate in the plasma. Utilizing the resist’s thicknesses before and after ICP-RIE, we compute the etch rates of the plasmas for the materials involved and determine selectivity ratios. Fig. 2 (b,c) summarizes the resulting, dose-dependent selectivities (details see caption). To estimate selectivity on diamond, we repeat evaluation for each ICP-RIE etched diamond using three randomly selected, but neighboring 25 m square structures for each diamond for the base dose for which we observed optimal writing quality. We note that even for these neighboring structures, we observe thickness variations. On diamond, we observe a decreased etch rate of the ”Pillar Etch” plasma on Medusa 84 SiH compared to silicon leading to increased selectivities for the chosen dose of 350 C/cm2. For a diamond featuring a Medusa 84 SiH layer of 130 nm 20 nm thickness, etched for 6 minutes, the selectivity was found to be 12 1. For a second diamond with higher layer thickness of 300 nm 10 nm, etched for 8 minutes we find a selectivity of 10.9 0.2. Both indicate that the insulating diamond substrate influences plasma dynamics and reduces the etch rate of the ”Pillar Etch” on Medusa 84 SiH by approximately 40 %, favoring its application in diamond nanofabrication. Moreover, beyond , there is a notable increase in selectivity. We suspect enhanced cross-linking and resist densification to cause this effect, indicating that further hard baking post-development could improve etch resistance at lower lithography dose. The selectivity values presented in this study are comparable to those previously reported for similar ICP-RIE recipes using HSQ or SiO2-like masks [21, 4].
@@ 3.4 Fabrication Yield
We analyze optical microscopy images of all diamond nanopillars fields to quantify the fabrication yield of nanopillars. We note that prior to this evaluation, we used SEM imaging to make sure the structures we observe in optical microscopy are intact nanopillars and not structures created due to collapsed pillar masks that adhered to their original positions. A total of 79.200 nanopillars have been evaluated on a single electronic grade diamond. We report a total fabrication yield of 89.6 %. Among the 10.4 % fabrication failures, 59.4 % result from a non-optimal synthesis and oligomeric silsesquioxanes [25], 20.7 % are caused by edge beads that affect lithography and 17.7 % are the result of adhesion failure. The remaining 2.2 % is due to contaminants, as well as overexposure in lithography. The work presented here is based on an early prototype of Medusa 84 SiH by Allresist. The now commercially available version has been significantly optimized regarding manufacturing to remove the effect originating from non-optimal synthesis. We note original FOx resist was also prone to similar issues before synthesis optimization [25]. Therefore fabrication yields could potentially reach around 96 %, primarily constrained by edge beads on small diamond substrates and sporadic adhesion failures.
@@ 4 Experimental Setup
We used a custom-built confocal microscope (modified version of [26]) to study the impact of diamond nanostructuring with Medusa 84 SiH on NV centers by comparing their room-temperature spin properties before and after processing. As our goal is to obtain pillars with a high probability of finding single NV, the NVs in the implanted layer are so dense, that we observe small ensembles in our confocal microscope instead of individual NV centers prior to nanopillar fabrication. We utilize a diode laser at 520 nm, featuring sub-nanosecond digital modulation (DL nSec, PE 520, Swabian Instruments) and a maximum output power of 40 mW to excite and initialize the NV centers. A 600 nm shortpass filter placed behind the single-mode fiber guiding the excitation light eliminates background light from the fiber. The microscope objective (LMPLFLN100X, Olympus) has a numerical aperture of 0.8, along with a working distance of 3.4 mm. We utilize a piezo-electric scanner (ANSxyz100, standard range, Attocube Systems) to facilitate sample scanning. We use two dichroic mirrors and a 600 nm longpass filter to separate excitation and fluorescence light. We sent the fluorescence via a single mode fiber to a fiber coupled avalanche photo diode (APD, SPCM-AQRH-14-FC, Excelitas Technologies) with a quantum efficiency of approximately 69 % at 700 nm. A data acquisition card (PCIe-6323, National Instruments) acquires and logs the signal. To analyze fluorescence spectra, we direct the light to a diffraction grating spectrometer (Acton SpectraPro SP-2500, Princeton Instruments) that records the spectra using a CCD camera (Pixis 256E, Princeton Instruments) using a multi-mode fiber replacing the single mode detection fiber. To investigate excited state lifetimes via time correlated single photon counting, we use a supercontinuum laser (SuperK Extreme EXW-12, NKT-Photonics) with a configurable single-line filter (SuperK VARIA, NKT-Photonics) and a time-tagging module (PicoHarp 300, PicoQuant). For coherent spin manipulation, a microwave source (TSG4100A, Tektronix) is connected to a +45 dB amplifier (ZHL-16W-43+, Mini-Circuits), feeding an-shaped stripline microwave antenna [27], on which the diamond is fixed using Crystal Bond 509 (Structure Probe Inc). We apply adjustable external magnetic fields using neodymium magnets on a rotation mount and a three-axis linear translation stage.
@@ 5 Shallow NV Centers in Nanopillars manufactured using Medusa 84 SiH Masks
Fig. 3 (a) illustrates a confocal xy-scan of a nanopillar array (diameters 214 nm 5 nm). The fluorescence intensity emitted from the nanopillars is typical for individual NV centers in such structures [30, 6, 7, 5, 4]. It should be noted that all the measurement data displayed graphically stem from nanopillars exhibiting single NV centers. This was checked by their brightness and the presence of maximally one resonance pair during optically detected magnetic resonance (ODMR) within differently oriented external magnetic fields. Fluorescence spectra of pillars reveal the spectrum of NV centers, characterized by their broad phononic sideband and a zero phonon line at 637 nm, as shown in the inset of Fig. 3 (b). We perform lifetime measurements and fit them using a bi-exponential decay convoluted with a Gaussian to account for the instrument response function of our setup (244 ps width) as shown in Fig. 3 (b). Using the bi-exponential decay is necessary as we observe a fast decaying fluorescence component that we attribute to (broadband) surface fluorescence ( 1.9 ns 0.1 ns) [28, 29] which does not hinder using these structures in quantum sensing. We determine the NV excited state lifetime to be 15 ns 2 ns, which is slightly increased in comparison to NV centers in bulk diamond due to their proximity to the surface and excludes excessive quenching. To estimate the number of NV centers per nanopillar, we evaluated xy-scans for nanopillar fields. By identifying the local maxima of the histogram in Fig. 3 (c) as occurring due to the presence of 0, 1, 2, or 3 NV centers and assuming Gaussian distributions, we approximated the occurrence of a single NV center per nanopillar as 18 % 7 % (details see caption) comparable to Ref. [6]. Using this distribution, we estimate the creation yield of shallow NV centers in our sample. We determined an NV center density of , indicating a creation yield of 1.7 % 0.8 % which is comparable to other works [6] excluding excessive charge-state instability. We note that this approach only roughly estimates the NV density as the expected photonic enhancement and thus the apparent brightness of NV centers in nanopillars has been found to strongly depend on the spatial position within the nanopillar (see, e.g. Ref. [7]). We note that the observed NV centers are stable and do not show blinking or bleaching. This observation together with the NV creation yield matching previous work indicate that Medusa 84 SiH efficiently protects shallow NV centers during ICP-RIE.
@@
We examine spin characteristics of the shallow NV ensemble before processing and for individual NV in nanopillars after processing to assess their usability as quantum sensors. To this end, we perform optically detected magnetic resonance in presence of an external magnetic field. After processing, we select nanopillars for further investigation with single NV centers. We select a single transition for resonant driving of Rabi oscillations, to determine the -pulse duration as shown in the inset of Fig. 3 (d). Subsequently, we perform spin echo experiments to determine the NV center’s coherence time , as illustrated in Fig. 4 (a). This procedure was repeated for 17 nanopillars, each containing a single NV center, we compare these values to of the shallow NV ensemble before nanopillar fabrication. We employ a fitting function from [26] to extract :
@@
In this study, we found an average coherence time of = 14 s 9 s before nanopillar fabrication using Medusa 84 SiH and = 9 s 6 s after processing. As shown in Fig. 4 (b), the coherence times are in the same window as observed before, which indicates that processing diamond using Medusa 84 SiH does not strongly reduce shallow NV center’s coherence times T2. The determined coherence times are in the same order of magnitude as those reported for single NV centers within nanopillars fabricated using FOx, where nanofabrication was also found to reduce the average [6]. Additionally, Ramsey measurements, as displayed in Fig. 4 (c), have been conducted. In average, we found dephasing times of = 1.2 s 0.4 s which lies in the same range as for commercially available single NV scanning probes [8]. Moreover, T1 spin lifetime measurements were performed before and after processing. As shown in Fig. 4 (d), the spin lifetime also shows only a slight decrease after processing (average T1 = 0.3 ms 0.1 ms ). While T1 before and after processing is consistent for our diamond sample, we note that we typically find T1 times of several milliseconds for commercially available NV scanning probes [8]. These results conclusively demonstrate that the use of Medusa 84 SiH for diamond nanofabrication, compared to the cutting-edge HSQ resist FOx by DuPont, does not adversely affect the spin properties of color centers, thus rendering Medusa 84 SiH appropriate for diamond-based quantum technologies.
@@ 6 Conclusion
This study examined Allresist’s novel electron beam resist Medusa 84 SiH as a substitute for DuPont’s Flowable Oxide for diamond nanofabrication, focusing on its use in diamond-based quantum technologies. We transferred previously established processes for FOx to Medusa 84 SiH [6, 7, 21] and proved their feasibility by realizing prototypical test structures at the m-scale down to nanopillars with diameters around 200 nm. We investigated the resist’s basic properties by evaluating fabrication yield, recording contrast curves and determining its dose-dependent etch resistance against common plasmas necessary to nanostructure diamond. We confirm that Medusa 84 SiH is very similar to FOx for the demonstrated prototypical structures with comparable contrast, achieved fabrication yields of up to 96 % for electron beam evaporated silicon as adhesion promoting interlayer and shows a conservatively estimated selectivity towards high biased O2 + Ar ICP-RIE plasma in diamond etching of 11 to 12 for m-scale structures. We further investigated the impact of processing diamond with Medusa 84 SiH on shallow NV centers within nanopillars as applied in the field of quantum sensing [6, 7, 8] by comparing T2, T and T1 times before and after fabricating nanopillars. We report negligible decreases in spin lifetime, dephasing and coherence times, therefore proving the suitability of these nanopillars for quantum sensing schemes. Further, we investigated their fluorescence emission and show that, with the proposed nanofabrication process, an enhancement in photon collection efficiency of 7.0 0.9 was achieved. Overall, Medusa 84 SiH is a promising alternative to FOx as an electron beam resist for diamond nanofabrication compatible with quantum technologies like NV sensing. Future optimization might include contrast optimization, e.g. lowering the exposure to thermal energy during soft baking and adjusting the development parameters. In addition to diamond, this work hints Medusa 84 SiH might be also an appropriate alternative for FOx in other material systems like silicon, SOI, GaAs or other materials under research requiring nanoscale resolution and high resistance towards RIE processes.
@@ 7.1 Nanofabrication
| Spin Coating of Resist Layer | Step #1: 1500 rpm for 6 s, Acc.: 500 rpm/s |
| Step #2: 4000 rpm for 30 s, Acc.: 1000 rpm/s | |
@@
| Application of conductive Layer | Resist: ESpacer 300Z, 4000 rpm for 30 s, Acc.: 1000 rpm/s |
@@
| Resist: Electra 92, 2000 rpm for 60 s, Acc.: 1000 rpm/s | |
| Layer Thickness: 60 nm, Soft Bake: 90 ∘C for 120 s | |
@@
| Rinse: Ultrapure water for 30 s, Cleaning: Isopropanol for 30 s | |
@@
Table 1 provides a comprehensive summary of all process steps in sequential order, detailing the majority of the process parameters specifically related to the electronic grade diamond. For diamonds of lower purity utilized in process development, the stress relief etch, ion implantation and high vacuum annealing steps were omitted. Moreover, the silicon samples used for contrast curve analysis did not undergo Tri-Acid or KOH treatments, and procedures such as electron beam evaporation, deposition of a conductive layer, and proximity correction were also excluded.
@@
Table 2 summarizes the parameters used to establish the ICP-RIE plasmas used to conduct stress relief etches on an Oxford Instruments Plasmalab 100 or to structure diamond with the proposed fabrication process using a Sentech PTSA-ICP Plasma Etcher SI 500. It should be noted that both ICP-RIE devices can realize the described processes, although their parameters require appropriate adjustment.